A race car moves such that its position fits the relationship

x = (5.5 m/s)t + (0.60 m/s3)t3 where x is measured in meters and t in seconds.

(a) A plot of the car's position versus time is which of the following?

x (m) 70 160 140 12 100 40 60 40 20 20 l0- t (s) t (s) 1.00 2.00 3.00 4.00 5.00 6.00 1.00 2.00 3.00 4.0 5.00 6.00 x (m) x (m) 180 160 140 12 100 20 15 5 40 20 t (s) 1.00 2.00 3.00 4.00 5.00 6.00 1.00 2.00 3.00 4.005.00 6.00

(b) Determine the instantaneous velocity of the car at t = 5.0 s, using time intervals of 0.40 s, 0.20 s, and 0.10 s. (In order to better see the limiting process keep at least three decimal places in your answer.)

Δt = 0.40 s

    

m/s (Use the interval from

t = 4.80 s to 5.20 s.)

Δt = 0.20 s

    
m/s (Use the interval from t = 4.90 s
to 5.10 s.)

Δt = 0.10 s

    
m/s (Use the interval from t = 4.95 s
to 5.05 s.)


(c) Compare the average velocity during the first 5.0 s with the results of part (b).

The average velocity of _______ m/s is (much less than, about the same as, much greater than) the instantaneous velocity.
Answer #1

To plot graph , assume that the values of t from 0 s to 6 s.by putting t values in the given equation

get corresponding x values .

d) Aho typical values are χ(n) om) t (G) ac (ro) 6.I 15.g 32.7 60.4 lo 2.5 160 yot Izo 20 62-6 え3us b) uoe voill use a oyos inteval centered at t ss. ㅈ = 112-96m 20 20Sm 50.5 m/s o 20s at t=y.qss, χこ99.9q구yn, akts5.oso, s os01m50.sm/s C) At t-56, χ= 102. 5m. Thus, bo He fin+ 5s ー-102-Sm-0-120.5 m/s This value i much less+ta the utantanecus velouty at t=53. 5S-0